OpenStudy (anonymous):

We define the sequence of numbers an = 1 if 0 <=n <=3 and an-1 + an-2 + an-3 + an-4 , if n >= 4. Prove that an =1 (mod 3) for all n >= 0.

OpenStudy (experimentx):

a6 = 11+2 = 13 = 1 mod 3 each term is 1 mod 3 ... so it's (1 + 1 + 1 + 1) mod 3 = 1 mod 3 ... something like that

OpenStudy (experimentx):